OpenStudy (anonymous):

Covert the function into standard form. y-2=-3(x+4)^2

OpenStudy (p0sitr0n):

develop and put everything except y to the right hand side y=-3(x^2+8x+16)+2 y=-3x^2-24x-48+2 y=-3x^2-24x-46 optionally you can arrange the exponents in decreasing order

OpenStudy (anonymous):

Thanks a lot!